y-2dy/dx + y-1= 2x


1
Expert's answer
2022-02-22T14:08:57-0500

The given DE can be written in the form:


"\frac{1}{y^2}\dfrac{dy}{dx} + \frac{1}{y}= 2x"

which is a Bernoulli type of first order DE.


"\text{Solve Bernoulli's equation }\frac{1}{y(x)}+\frac{\frac{d y(x)}{d x}}{y(x)^{2}}=2 x:\\[3mm]
\text{Multiply both sides by }-1 :\\
-\frac{\frac{d y(x)}{d x}}{y(x)^{2}}-\frac{1}{y(x)}=-2 x\\
\text{Let }v(x)=\frac{1}{y(x)}, \text{ which gives }\frac{d v(x)}{d x}=-\frac{\frac{d y(x)}{d x}}{y(x)^{2}}:\\
\frac{d v(x)}{d x}-v(x)=-2 x \text{ Let }\mu(x)=e^{\int-1 d x}=e^{-x}.\\[3mm]
\text{Multiply both sides by }\mu(x) :\\
e^{-x} \frac{d v(x)}{d x}-e^{-x} v(x)=-2 e^{-x} x\\
\text{Substitute }-e^{-x}=\frac{d}{d x}\left(e^{-x}\right):\\
e^{-x} \frac{d v(x)}{d x}+\frac{d}{d x}\left(e^{-x}\right) v(x)=-2 e^{-x} x\\"


"\text{Apply the reverse product rule } f \frac{d g}{d x}+g \frac{d f}{d x}=\frac{d}{d x}(f g)\text{ to the left-hand side}:\\
\frac{d}{d x}\left(e^{-x} v(x)\right)=-2 e^{-x} x\\[3mm]
\text{Integrate both sides with respect to }x :\\[3mm]
\int \frac{d}{d x}\left(e^{-x} v(x)\right) d x=\int-2 e^{-x} x d x\\[3mm]

\text{Evaluate the integrals}:\\
e^{-x} v(x)=-2 e^{-x}(-x-1)+c_{1},\text{ where } c_{1} \text{ is an arbitrary constant.}\\[3mm]
\text{Divide both sides by }\mu(x)=e^{-x}:\\[3mm]

v(x)=2 x+c_{1} e^{x}+2"

Solving for "y(x)", we have:


"y(x)=\frac{1}{v(x)}=\frac{1}{2 x+c_{1} e^{x}+2}"



Which is the required solution of the given DE.